OpenStudy (anonymous):

Can someone very simple explain this and use an example? Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (anonymous):

This question i dont quite understand? Please someone help?

OpenStudy (anonymous):

im assuming by two similar figure they mean the same figure with different dimensions, so use a figure like a cube that is easy to solve for surface area and volume

OpenStudy (anonymous):

Ok and then?

OpenStudy (anonymous):

Like can you show me how to do these problems? Its the last question of the night and im stumped ._.

OpenStudy (anonymous):

so say the first cube has length=width=height=2 and the second l=w=h=4 solve for the surface area and the volume of each to start

OpenStudy (anonymous):

Wait so they all equal 2? Then its be 8? Wouldnt it?

OpenStudy (anonymous):

And the second one would be 64?

OpenStudy (anonymous):

volume=l*w*h surface area = 6*l*w for cubes

OpenStudy (anonymous):

Bleh im sorry ._. its late and im like braindead.

OpenStudy (anonymous):

the similarity ratio for these would be 1/2 because the first cubes dimensions are 1/2 of the second cubes

OpenStudy (anonymous):

ive been there dont worry

OpenStudy (anonymous):

nnow we need to solve for the volumes and surface areas with the equations i gave you and find their ratios

OpenStudy (anonymous):

Alright, so we have to solve there two numbers you gaves me for volume and surface areas using there formulas?

OpenStudy (anonymous):

sorry ratio of surface area to volume would be SA/V for each cubes (the ratio will be the same for both)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So i do what i just said?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

But you only gave me 2 numbers. when i need more then that? You gave me 2 and 4.

OpenStudy (anonymous):

so for the first cube length=2in height=2in and width =2in so volume is 2in*2in*2in=8in^3 surface area =6in*2in*2in=24in^2 i just decided to use inches for my units but it could be any unit of distance

OpenStudy (anonymous):

sorry surface area =6*2in*2in

OpenStudy (anonymous):

Whered you get the 6 from?

OpenStudy (anonymous):

because the area of a square is its hieght times its width and there are six 'faces' of a cube that are all squares

OpenStudy (anonymous):

so its surface area is six times the area of on of those squares

OpenStudy (anonymous):

Ohhhh.

OpenStudy (anonymous):

now the ratio of surface area to volume would be 24/8=3

OpenStudy (anonymous):

if you solve for the surface area and volume of the cube whose length width and height all equal four you will find its ratio to also be 3

OpenStudy (anonymous):

Snap your a genius :D

OpenStudy (anonymous):

Alright, so to answer my 3 questions basically, 1. (2,2,2) (6,2,2) 2. The similiarity rate is 24/8=3 And 3 is just where i show how i did it?

OpenStudy (anonymous):

NOOOOOOOOOOOOOOO, you went offline ._.

OpenStudy (anonymous):

Can someone please finish this? Did i do what i need to submit correctly? And if not could you please help me?

OpenStudy (anonymous):

sure